# The relevance of degenerate states in chiral polaritonics

### v1

@@ Abstract
In this work we explore theoretically whether a parity-violating/chiral light-matter interaction is required to capture all relevant aspects of chiral polaritonics or if an parity-conserving/achiral theory is sufficient (e.g. long-wavelength/dipole approximation). This question is non-trivial to answer, since achiral theories (Hamiltonians) still possess chiral solutions. To elucidate this fundamental theoretical question, a simple GaAs quantum ring model is coupled to an effective chiral mode of a single-handedness optical cavity in dipole approximation. The bare matter GaAs quantum ring possesses a non-degenerate ground state and a doubly degenerate first excited state. The chiral or achiral nature (superpositions) of the degenerate excited state remains undetermined for an isolated matter system. However, inside our parity-conserving description of a chiral cavity, we find that the dressed eigenstates automatically (ab-initio) attain chiral character and become energetically discriminated based on the handedness of the cavity. In contrast, the non-degenerate bare matter state (ground state) does not show an energetic discrimination inside a chiral cavity within dipole approximation. Nevertheless, our results suggest that the handedness of the cavity can still be imprinted onto these states (e.g. angular momentum and chiral current densities). Overall, above findings highlight the relevance of degenerate states in chiral polaritonics. In particularly because recent theoretical results indicate the formation of a frustrated and highly-degenerate electronic ground-state under collective strong coupling conditions, which thus might be prone to chiral symmetry breaking effects.
@@ I Introduction
Irrespective of these theoretical subtleties, matter systems that exhibit chirality are crucial in many fields of natural science [5]. Specifically in chemistry, molecules that exhibit chirality (enantiomers) are of central importance. While enantiomers have similar chemical and physical properties, they react differently when interacting with other chiral objects. Hence controlling and differentiating between enantiomers is important in chemical synthesis. To do so it is common to use chiral solvents or substrates, but this practice is often expensive and, in many cases, not environmentally friendly [6, 7]. An alternative is to use chiral optical fields. For instance, circularly polarized light is used for detecting chiral substances [8, 9, 10] and theoretical as well as experimental works show its potential to control chemical processes enantioselectively [11, 12, 13]. More recently the idea of using optical cavities with chiral modes (chiral cavities) has been considered a promising tool for enantiomeric discrimination and manipulation [14, 15, 16, 17, 18, 19, 20, 21, 22]. Of specific interest is the case of collective strong coupling [23, 24, 25, 26], where a large ensemble molecules is coupled to the modes of a cavity.
@@
As can be anticipated from the previous theoretical ambiguities, there is no straightforward and unified approach to chose a Hamiltonian that captures all chemically relevant aspects of a molecular ensemble that couples strongly to a chiral cavity. A specific theoretical aspect that we want to investigate in this work is whether we need to use an explicit parity-violating light-matter coupling term in the Hamiltonian or whether certain chiral properties can also be captured due to solutions with parity-conserving light-matter coupling terms. In particular, choosing a parity-conserving light-matter coupling term may be computational beneficial when aiming for collective strong-coupling simulations with large molecular ensembles. The prime example of such a computationally efficient parity-conserving coupling term, even for chiral cavity modes, is the long-wavelength/dipole light-matter coupling [27]. The obvious advantage of a parity-violating light-matter coupling term for polaritonics is that it allows to energetically discriminate between different enantiomers in a chiral cavity even for non-degenerate states. If, however, the light-matter coupling term is parity conserving, i.e., the interaction with a chiral cavity transforms back onto itself under parity, then both enantiomers will have the same energy in the chiral cavity. Nevertheless, this only tells that there is no energetic discrimination for non-degenerate states, but still the wave function or other observables may have chiral character. Furthermore, since we are usually interested in the collective-coupling case, which implies highly degenerate ensemble states [28], cavity-induced spontaneous symmetry breaking becomes plausible. This in turn can affect the chiral properties of the solutions, and might become the dominating effect for chiral polaritonics. In other words, even a parity-conserving coupling term might be enough to capture the essentials of chiral polaritonics in the collective case.
@@
Taking a step towards this direction, in this work we consider a simple toy system and investigate to which extend we need a parity-violating light-matter coupling term to describe/control chiral properties. We do so for an achiral matter system that is coupled to a single-handed chiral cavity field, [17, 18] within the local strong coupling regime. That is, we do not consider a large ensemble that couples collectively to the cavity mode, but instead a single entity that couples strongly. To be specific, we consider a two-dimensional quantum ring that couples to an effective single-mode cavity. This can be taken as a model for a single quantum ring coupled strongly to a micro- or nano-cavity or a single quantum ring out of an ensemble that locally couples strongly due to inherent disorder in the ensemble [29, 28]. We will investigate how various observables behave and to which extend they differ between a parity-violating and parity-conserving light-matter coupling. In the end we will discuss to which extend these results can help to anticipate what happens for more realistic setups and in the collective-coupling case of chiral polaritonics.
@@ II Theory and methodology
Thus we have quantized the photon field in a chiral representation [...]. The connection to the linear-polarized quantization is given by expressing
@@
for $k>0$. See Fig. 1 for a artistic sketch of the total setup.
Figure 1: Artistic sketch of the setup. We consider a single effective electron that is restricted to two dimensions in an external ”Mexican-hat”-like potential. The effective electron interacts with a quantized chiral photon mode polarized along the $z$-axis. The different handedness of the cavity is represented by their corresponding spiral orientation.
@@
Let us next consider a simplified description of a chiral cavity. Here we focus on the long-wavelength/dipole approximation. That is, we assume that the wavelength of the cavity-enhanced modes is small compared to the spatial extend of the matter system such that we can approximate $e^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}\approx 1$. This approximation introduces extra symmetries. On the one hand, for each mode in Eq. (2) we have now
@@
In order to now compare an explicitly parity-violating and parity-non-violating description of our chiral-cavity setup, we in the following use a single circularly-polarized mode (with either handedness + or - as shown in Fig. 1) to mimic a single-handedness[17, 18] chiral cavity. and for the explicitly parity-violating interaction we include the external magnetic field of Eq. (8). We will compare to which extend these two descriptions of a chiral-cavity setup differ and to which extend the parity-conserving description of just a circularly-polarized mode in dipole approximation can capture/control the chiral character of the coupled wave functions and their observables.
For our calculations, unless otherwise stated, we use a frequency of of the chiral cavity mode $\hbar\omega=1.413$ meV, which is in resonance with the first transition of the uncoupled matter system. To simulate the system, we used a two-dimensional grid of 71×71 points, with a separation of 1.27 nm between each pair of points, and 9 photon states.
@@ III.1 Uncoupled matter system
Figure 4: Magnitude of the transition terms, in atomic units, for A) the $x$ or $y$ components of the dipole moment, B) the $z$ component of the magnetic dipole moment, C) the $xx$ or $yy$ components of the quadrupole tensor, and D) the $xy$ component of the quadrupole tensor, between the ground state (gs) and the states $M_{\pm}$, which correspond to the degenerate first excited state of the uncoupled system. The main contributions are highlighted in red.
@@ III.2 Chiral cavity without magnetic field
Next we consider how the bare matter picture changes upon coupling to a quantized effective (single) chiral cavity mode in the long-wavelength approximation. As discussed in Sec. II, the chosen chiral light-matter interaction will not break parity explicitly. It turns out that the cavity cannot couple equally to the two chiral eigenstates of the bare matter problem $|M_{\pm}\rangle$. In particular, the handedness of the cavity breaks the degeneracy of the excited states by coupling mostly with one state and creating two polaritons. Thus a chiral cavity in dipole approximation couples selectively to the chiral superpositions (solutions) obtained from an the achiral bare matter theory (Hamiltonian). Notably, this chiral discriminating nature of the excited state couplings emerges naturally (ab-initio), whereas for the bare matter system the chiral nature only appears upon a specific choice (superposition) of the degenerate excited states. In other words, no ad hoc assumption was necessary, i.e., no a priori restriction to a chiral basis of the bare matter system was imposed as, e.g., done in Ref. [15]. Notice, however, the handedness of the energy spectrum becomes only interpretable / assignable if the handedness of a single-handed cavity is known.
@@
In Fig. 5 we present the dipole spectra for different coupling strengths. For our case we change $\lambda=\sqrt{\hbar|e|^{2}/\epsilon_{0}L^{3}}$ by modifying $|e|^{2}/(\epsilon_{0}L^{3})$ while keeping the frequency $\omega$ the same. In all our results, we express $\lambda$ in atomic units. Each spectrum obtained in the cavity shows three peaks: the first and third peaks correspond to the polaritonic states, and the second peak corresponds to the state that weakly couples with the cavity, which we thus label according to the bare matter nomenclature as $M_{\pm}$ (the sign is opposite to the one of the cavity). The obtained Rabi splitting follows the usual linear dependency with respect to the coupling parameter $\lambda$. However, the separation of the peaks is not symmetric with respect to the cavity frequency. Similarly the peak assigned to the weakly coupled state is red shifted as we increase the coupling strength. These changes are attributed to the “detuned” interaction between states and the back-action of the matter system onto the cavity.[28, 33, 34, 35] We will come back to this in the next section.
@@
In Fig. 6 we can see the changes in $L_{z}$ of the ground state and the first three excited states. For lower coupling strength, the two polaritons share half of the angular momentum of the original state $|M_{\pm}\rangle$, while the uncoupled state is not considerably affected. As the coupling strength increases, $L_{z}$ of the upper polariton (UP) and the lower polariton (LP) separates, and the last one gets closer to the one of the ground state which moves away from zero. Thus also the ground state gets a considerable chiral character (angular currents). The value of $|L_{z}|$ decreases for the weakly-coupled state with resepct to the coupling strength. This is because, at higher coupling strengths, the cavity imprint its opposite angular momentum. The same happens with the ground state. Furthermore, notice that the rigorous bare matter angular selection rule are weakened by the cavity for electronic dipole transitions. Now, the transitions from the ground state to any of the first excited states involves $10<|\Delta M|<$ 1, at least within the numerically explored parameter range of $\lambda$. The presence of a chiral cavity in dipole approximation also enables transitions between the polaritonic states through magnetic-dipole coupling (see Fig. 7), which is absent for the bare system.
Figure 7: A) Magnetic-dipole moment ($m_{z}$) matrix obtained for $\lambda$ = $5\times 10^{-6}$ without external magnetic field. Relevant transitions are highlighted in red. B) Magnetic-dipole moment transition between LP and UP are plotted with respect to different coupling strengths. In both plots, the magnetic-dipole moment is expressed in atomic units. The results are independent of the cavity polarization.
@@
Figure 8: Matrices for the A) xx (or yy) and B) xy components of the quadrupolar tensor, obtained for the first four states, using $\lambda=5\times 10^{-6}$, in atomic units. The shown magnitudes are independent of the cavity handedness. The main contributions are highlighted in red. The cavity-induced transitions do not follow the original selection rule $\Delta M=\pm 2$.
@@ III.3 Chiral cavity with magnetic field
Next we want to see whether qualitative changes appear in the coupled light-matter wave functions and observables, once we employ an explicit parity-violating interaction term. In particular, we expect that the non-degenerate bare matter ground-state is energetically discriminated by the chiral photon field. The reason is that the parity-violating part of the photon field imprints a handedness onto the non-degenerate bare-matter ground-state. It thus couples differently to the two circular polarizations of the cavity. In our simulations $B_{0}=0.235$ T was used. A comparison of the coupled ground-state energy discrimination in a parity-conserving chiral cavity with and without external magnetic fields is given in Fig. 9.

### v2

@@ Abstract
In this work we explore theoretically whether a parity-violating/chiral light-matter interaction is required to capture all relevant aspects of chiral polaritonics or if a parity-conserving/achiral theory is sufficient (e.g. long-wavelength/dipole approximation). This question is non-trivial to answer, since achiral theories (Hamiltonians) still possess chiral solutions. To elucidate this fundamental theoretical question, a simple GaAs quantum ring model is coupled to an effective chiral mode of a single-handedness optical cavity in dipole approximation. The bare matter GaAs quantum ring possesses a non-degenerate ground state and a doubly degenerate first excited state. The chiral or achiral nature (superpositions) of the degenerate excited states remains undetermined for an isolated matter system. However, inside our parity-conserving description of a chiral cavity, we find that the dressed eigenstates automatically (ab-initio) attain chiral character and become energetically discriminated based on the handedness of the cavity. In contrast, the non-degenerate bare matter state (ground state) does not show an energetic discrimination inside a chiral cavity within dipole approximation. Nevertheless, our results suggest that the handedness of the cavity can still be imprinted onto these states (e.g. angular momentum and chiral current densities). Overall, above findings highlight the relevance of degenerate states in chiral polaritonics. In particular, because recent theoretical results for linearly polarized cavities indicate the formation of a frustrated and highly-degenerate electronic ground-state under collective strong coupling conditions, which, likewise, is expected to form in chiral polaritonics and thus could be prone to chiral symmetry breaking effects.
@@ I Introduction
Irrespective of these theoretical subtleties, matter systems that exhibit chirality are crucial in many fields of natural science [5]. Specifically in chemistry, molecules that exhibit chirality (enantiomers) are of central importance. While enantiomers have similar chemical and physical properties, they react differently when interacting with other chiral objects. Hence controlling and differentiating between enantiomers is important in chemical synthesis. To do so it is common to use chiral solvents or substrates, but this practice is often expensive and, in many cases, not environmentally friendly [6, 7]. An alternative is to use chiral optical fields. For instance, circularly polarized light is used for detecting chiral substances [8, 9, 10] and theoretical as well as experimental works show its potential to control chemical processes enantioselectively [11, 12, 13]. More recently the idea of using optical cavities with chiral modes (chiral cavities) has been considered a promising tool for enantiomeric discrimination and manipulation [14, 15, 16, 17, 18, 19, 20, 21, 22]. Of specific interest is the case of collective strong coupling [23, 24, 25, 26], where a large ensemble of molecules is coupled to the modes of a cavity.
@@
As can be anticipated from the previous theoretical ambiguities, there is no straightforward and unified approach to chose a Hamiltonian that captures all chemically relevant aspects of a molecular ensemble that couples strongly to a chiral cavity. A specific theoretical aspect that we want to investigate in this work is whether we need to use an explicit parity-violating light-matter coupling term in the Hamiltonian or whether certain chiral properties can also be captured due to solutions with parity-conserving light-matter coupling terms. In particular, choosing a parity-conserving light-matter coupling term may be computationally beneficial when aiming for collective strong-coupling simulations with large molecular ensembles. The prime example of such a computationally efficient parity-conserving coupling term, even for chiral cavity modes, is the long-wavelength/dipole light-matter coupling [27]. The obvious advantage of a parity-violating light-matter coupling term for polaritonics is that it allows to energetically discriminate between different enantiomers in a chiral cavity even for non-degenerate states. If, however, the light-matter coupling term is parity conserving, i.e., the interaction with a chiral cavity transforms back onto itself under parity, then both enantiomers will have the same energy in the chiral cavity. Nevertheless, this only tells that there is no energetic discrimination for non-degenerate states, but still the wave function or other observables may have chiral character. Furthermore, since we are usually interested in the collective-coupling case, which implies highly degenerate ensemble states [28], cavity-induced spontaneous symmetry breaking becomes plausible. This in turn can affect the chiral properties of the solutions, and might become the dominating effect for chiral polaritonics. In other words, even a parity-conserving coupling term might be enough to capture the essentials of chiral polaritonics in the collective case.
@@
Taking a step towards this direction, in this work we consider a simple toy system and investigate to which extent we need a parity-violating light-matter coupling term to describe/control chiral properties. We do so for an achiral matter system that is coupled to a single-handed chiral cavity field, [17, 18] within the local strong coupling regime. That is, we do not consider a large ensemble that couples collectively to the cavity mode, but instead a single entity that couples strongly. To be specific, we consider a two-dimensional quantum ring that couples to an effective single-mode cavity. This can be taken as a model for a single quantum ring coupled strongly to a micro- or nano-cavity or a single quantum ring out of an ensemble under collective strong coupling conditions. In the later case, it has been shown numerically that disorder can induce non-negligible local effects [29, 28] that one can mimic by computationally accessible single entity strong coupling simulations . We will investigate how various observables behave and to which extent they differ between a parity-violating and parity-conserving light-matter coupling. In the end we will discuss to which extent these results can help to anticipate what happens for more realistic setups and in the collective-coupling case of chiral polaritonics.
@@ II Theory and methodology
Thus we have quantized the photon field in a chiral representation. The connection to the linear-polarized quantization is given by expressing
@@
for $k>0$. In Fig. 1 the bare matter potential energy surface is depicted, the corresponding electronic ground state density of the quantum ring as well as a scheme of the total setup including the cavity.
Figure 1: We consider a single effective electron that is restricted to two dimensions in an external ”Mexican-hat”-like potential (top left), which gives rice to a quantum ring electronic density (bottom left). The effective electron interacts with a quantized chiral photon mode polarized along the $z$-axis (right). The different handedness of the cavity is represented by their corresponding spiral orientation.
@@
Let us next consider a simplified description of a chiral cavity. Here we focus on the long-wavelength/dipole approximation. That is, we assume that the wavelength of the cavity-enhanced modes is small compared to the spatial extent of the matter system such that we can approximate $e^{\mathrm{i}\mathbf{k}\cdot\mathbf{r}}\approx 1$. This approximation introduces extra symmetries. On the one hand, for each mode in Eq. (2) we have now
@@
In order to now compare an explicitly parity-violating and parity-non-violating description of our chiral-cavity setup, we in the following use a single circularly-polarized mode (with either handedness + or - as shown in Fig. 1) to mimic a single-handedness[17, 18] chiral cavity. and for the explicitly parity-violating interaction we include the external magnetic field of Eq. (8). We will compare to which extent these two descriptions of a chiral-cavity setup differ and to which extent the parity-conserving description of just a circularly-polarized mode in dipole approximation can capture/control the chiral character of the coupled wave functions and their observables.
For our calculations, unless otherwise stated, we use a frequency of of the chiral cavity mode $\hbar\omega=1.413$ meV, which is in resonance with the first transition of the uncoupled matter system. To simulate the system, we worked in a two-dimensional real-space grid of 71×71 points, with a separation of 1.27 nm between each pair of points, and 9 photon states. The chosen grid separation provides converged energies at sufficient computational speed. The energy and wave functions of each state were solved by numerically exact diagonalization of the full Hamiltonian operator. All the observables presented in this work where calculated by numerical integration of the corresponding exact eigenstates.
@@ III.1 Uncoupled matter system
Figure 4: Magnitude of the transition terms, in atomic units for the bare system, for A) the $x$ or $y$ components of the dipole moment, B) the $z$ component of the magnetic dipole moment, C) the $xx$ or $yy$ components of the quadrupole tensor, and D) the $xy$ component of the quadrupole tensor, between the ground state (gs) and the states $M_{\pm}$. The main contributions are highlighted in red.
@@ III.2 Chiral cavity without magnetic field
Next we consider how the bare matter picture changes upon coupling to a quantized effective (single) chiral cavity mode in the long-wavelength approximation. The corresponding Hamiltonian is given by,
$\hat{H}_{\pm}=\frac{1}{2m}\left(-\mathrm{i}\hbar\boldsymbol{\nabla}+\frac{\lambda}{\sqrt{2\omega}}\left(\boldsymbol{\epsilon}_{\pm}\hat{b}+\boldsymbol{\epsilon}_{\pm}^{*}\hat{b}^{\dagger}\right)\right)^{2}+V_{\rm ext}(\mathbf{r})+\hbar\omega\left(\hat{b}^{\dagger}\hat{b}+\frac{1}{2}\right),$ (13)
where $\pm$ labels the handedness of the cavity. The coupling constant between light and matter is defined as $\lambda=\sqrt{\hbar|e|^{2}/\epsilon_{0}L^{3}}$, where $L^{3}$ corresponds to the mode volume of the cavity. As discussed in Sec. II, the chosen chiral light-matter interaction will not break parity explicitly. It turns out that the cavity cannot couple equally to the two chiral eigenstates of the bare matter problem $|M_{\pm}\rangle$. In particular, the handedness of the cavity breaks the degeneracy of the excited states by coupling mostly with one state and creating two polaritons. Thus a chiral cavity in dipole approximation couples selectively to the chiral superpositions (solutions) obtained from the achiral bare matter theory (Hamiltonian). Notably, this chiral discriminating nature of the excited state couplings emerges naturally (ab-initio), whereas for the bare matter system the chiral nature only appears upon a specific choice (superposition) of the degenerate excited states. In other words, no ad hoc assumption was necessary, i.e., no a priori restriction to a chiral basis of the bare matter system was imposed as, e.g., done in Ref. [15] (see numerical justification below). Notice, however, the handedness of the energy spectrum becomes only interpretable / assignable if the handedness of a single-handed cavity is known.
@@
In Fig. 5 we present the dipole spectra for different coupling strengths ($\lambda$), by modifying $|e|^{2}/(\epsilon_{0}L^{3})$ while keeping the frequency $\omega$ the same. In all our results, we express $\lambda$ in atomic units. Each spectrum obtained in the cavity shows three peaks: the first and third peaks correspond to the polaritonic states, and the second peak corresponds to the state $|\Psi_{m(\pm)}\rangle$ that only weakly couples with the cavity by solving Eq. (13) numerically exactly. In more detail, investigating the projection of the bare matter chiral-superposition states $M_{\pm}$, defined in Eq. (12), reveals $|\langle M_{\pm}|\Psi_{m(\pm)}\rangle|^{2}\in[1,0.96)$ for the investigated coupling strengths $\lambda$ of a cavity described by $\hat{H}_{\mp}$. Due to this large overlap, we consider the middle state $|\Psi_{m(\pm)}\rangle$ as virtually unaffected by the cavity and label it, from now on, as $M_{\pm}$ (the sign is opposite to the one of the cavity). From this argument, we have justified ab-initio that a chiral cavity couples virtually only to the bare matter states that possess the same handedness as the cavity. The resulting Rabi splitting obeys the usual linear dependency with respect to the coupling parameter $\lambda$. However, the separation of the peaks is not symmetric with respect to the cavity frequency. Similarly the peak assigned to the weakly coupled state is red shifted as we increase the coupling strength. These changes are attributed to the “detuned” interaction between states and the back-action of the matter system onto the cavity.[28, 33, 34, 35] We will come back to this in the next section.
@@
In Fig. 6 we can see the changes in $L_{z}$ of the ground state and the first three excited states. For lower coupling strength, the two polaritons share half of the angular momentum of the original state $|M_{\pm}\rangle$, while the weakly-coupled state is not considerably affected. As the coupling strength increases, $L_{z}$ of the upper polariton (UP) and the lower polariton (LP) separates, and the last one gets closer to the one of the ground state which moves away from zero. Thus also the ground state gets a considerable chiral character (angular currents). The value of $|L_{z}|$ decreases for the weakly-coupled state with respect to the coupling strength. This is because, at higher coupling strengths, the cavity imprint its opposite angular momentum. The same happens with the ground state. Furthermore, notice that the rigorous bare matter angular selection rule are weakened by the cavity for electronic dipole transitions. Now, the transitions from the ground state to any of the first excited states involves $10<|\Delta M|<$ 1, at least within the numerically explored parameter range of $\lambda$. The presence of a chiral cavity in dipole approximation also enables transitions between the polaritonic states through magnetic-dipole coupling (see Fig. 7), which is absent for the bare system.
Figure 7: A) Magnetic-dipole moment ($m_{z}$) matrix for a cavity-coupling of $\lambda$ = $5\times 10^{-6}$ without additional external magnetic field. Relevant transitions are highlighted in red. B) Magnetic-dipole moment transition between LP and UP are plotted with respect to different coupling strengths. In both plots, the magnetic-dipole moment is expressed in atomic units. The results are independent of the cavity polarization.
@@
Figure 8: Matrices for the A) xx (or yy) and B) xy components of the quadrupolar tensor, obtained for the first four states inside a cavity, using $\lambda=5\times 10^{-6}$, in atomic units. No external magnetic field is applied. The shown magnitudes are independent of the cavity handedness. The main contributions are highlighted in red. The cavity-induced transitions do not follow the original selection rule $\Delta M=\pm 2$.
@@ III.3 Chiral cavity with magnetic field
Next we want to see whether qualitative changes appear in the coupled light-matter wave functions and observables, once we employ an explicit parity-violating interaction term. For this purpose, we apply a static external magnetic field $B_{0}$ in $z$-direction. The corresponding Hamiltonian is then given
$\hat{H}_{\pm}=\frac{1}{2m}\left(-\mathrm{i}\hbar\boldsymbol{\nabla}+\frac{\lambda}{\sqrt{2\omega}}\left(\boldsymbol{\epsilon}_{\pm}\hat{b}+\boldsymbol{\epsilon}_{\pm}^{*}\hat{b}^{\dagger}\right)-\frac{|e|B_{0}}{2}\left(y\mathbf{e}_{x}-x\mathbf{e}_{y}\right)\right)^{2}+V_{\rm ext}(\mathbf{r})+\hbar\omega\left(\hat{b}^{\dagger}\hat{b}+\frac{1}{2}\right).$ (15)
Explicitly adding a parity-violating interaction term, we expect that the non-degenerate bare matter ground-state is energetically discriminated by the chiral photon field. The reason is that the parity-violating part of the photon field imprints a handedness onto the non-degenerate bare-matter ground-state. It thus couples differently to the two circular polarizations of the cavity. In our simulations $B_{0}=0.235$ T was used. A comparison of the coupled ground-state energy discrimination in a parity-conserving chiral cavity with and without external magnetic fields is given in Fig. 9.